OpenStudy (anonymous):

Calculus II question. Evaluating Integral with trig functions... question posted inside post

OpenStudy (turingtest):

\[\int\limits_{\sqrt{\pi/2}}^{\sqrt{\pi}} 7\theta^3 \cos (\theta^2) d \theta \]?

OpenStudy (anonymous):

\[ \int\limits\limits_{\sqrt{\pi/2}}^{\sqrt{\pi}} 7\theta^3 \cos (\theta^2) d \theta \]

OpenStudy (turingtest):

I think the easiest way is to start with integration by parts

OpenStudy (anonymous):

yes. sorry! I am not fast at the equation thing

OpenStudy (turingtest):

lol it takes practice :)

myininaya (myininaya):

substitution could be useful

OpenStudy (anonymous):

i would think t=theta^2 dt= 2theta dtheta

OpenStudy (anonymous):

or 1/2dt= thetadtheta

OpenStudy (anonymous):

idk tho

OpenStudy (turingtest):

looks good to me so far

OpenStudy (anonymous):

give me a min to write and see what i got plz i am not fast

OpenStudy (anonymous):

\[so then I would have \int\limits_{\sqrt{\pi/2}}^{\sqrt{\pi}} 7\theta^3 \cos \theta^2 \theta d \theta\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\int\limits_{\sqrt{\pi/2}}^{\sqrt{\pi}} \cos t 1/2 dt\]

OpenStudy (anonymous):

meant to throw a 7 theta ^3 before cos :(

myininaya (myininaya):

\[\int\limits_{\sqrt{\frac{\pi}{2}}}^{\sqrt{\pi}}7 \theta \theta^2 \cos(\theta^2) d \theta\] so you mean to replace the theta^2 with t and the theta d theta with 1/2 dt so you get this: \[\int\limits_{\frac{\pi}{2}}^{\pi} 7 t \cos(t) \frac{1}{2} dt \]

myininaya (myininaya):

i also changed the limits

OpenStudy (anonymous):

ok.. do i plug into the integral udv = vu- integral vdu now?

OpenStudy (anonymous):

or i make the dv v u and du 1st and then plug in right?

OpenStudy (turingtest):

you need to decide on which is your u and which is your dv, yes

OpenStudy (anonymous):

would dv be 7 cos t dt ?

myininaya (myininaya):

sounds fine

myininaya (myininaya):

i guess you will do 1/2 t for the differentiating part

OpenStudy (anonymous):

for u?

myininaya (myininaya):

yah

OpenStudy (anonymous):

so then

myininaya (myininaya):

it doesn't matter where you put the constant though like you could leave the constant off until the end

OpenStudy (anonymous):

dv = 7 cos t dt v= 7 sin t u = 1/2 t du= 1/2 dt ?

OpenStudy (anonymous):

du doesnt look right. what did i do wrong?

OpenStudy (anonymous):

im sorry... i really need step by step help with this one! i will fan and reward promise

OpenStudy (turingtest):

you didn't substitute t for theta right myin corrected you but you didn't notice, go back up and look at the substitution she made

OpenStudy (anonymous):

i am lost :( already

OpenStudy (turingtest):

you had\[t=\theta^2\implies\theta d\theta=\frac12dt\]and your integrand was\[\theta^3\cos(\theta^2)d\theta=\theta^2\cos(\theta^2)\theta d\theta\]try substituting for t again; a littlemore carefully this time

OpenStudy (anonymous):

i will start completely over on this question. plz give me a min and hang in there with me

OpenStudy (anonymous):

what do i do with the 7 at this point?

OpenStudy (anonymous):

nothing?

OpenStudy (turingtest):

the 7 can be left out and then factored back in at the end, because\[\int Cf(x)dx=C\int f(x)dx\]if \(C\) is a constant. so you can asically just do \[\int \theta^3\cos(\theta^2)d\theta\]and multiply everything by 7 at the end

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

basically* Sometimes you actually want to bring the constant into the integration, because it makes the substitution simpler, but this is not one of those times.

OpenStudy (anonymous):

this question is very difficult

OpenStudy (turingtest):

just go step-by-step what was your substitution?

OpenStudy (anonymous):

if t= 1/2dt and dt= theta^2 cos theta^2 theta d theta

OpenStudy (anonymous):

7 t cos(t) dt

OpenStudy (turingtest):

good sorry I was afk

OpenStudy (turingtest):

yes, now integrate that by parts\[u=7t\\dv=\cos tdt\]